# Deep learning architecture for decrypting information on the event horizon

### v7

@@ 3 Deep learning for holographic system
Here, white represents a living cell and black represents a dead cell. During the inference process, the well-known glider pattern from the game of life was used as the input. However, it can be confirmed that the output after applying is the same as the output after applying , even with different inputs. Therefore, the proposed algorithm can learn a function that decrypts information pieces on a toy model of the event horizon to information pieces on a toy model of the conformal boundary using the commutativity. Because deep learning is a non-linear transform, this model can include non-trivial interactions if more physical constraints are added to the proposed model.
4 Discussion
In the proposed model, it was essential to assume that a time evolution law exists for many-body system, as in cellular automata. Indeed, this rule is not as simple as block CA and can not be expressed in a mathematically closed form. Therefore, if this rule exists, how to find it is a difficult problem. But again, we can expect that this rule can also be found using deep learning algorithm. First of all, we have a way to test a system in the standard model. Therefore, if datasets are collected by conducting numerous experiments, the evolution law will be able to be learned. In fact, if we assume that random data really evolves according to the block CA rule as an example in the previous section, it has been confirmed that the block CA rule can be learned through an appropriate deep learning architecture and numerous random dataset. Moreover, it can be confirmed that learning is possible in reverse time as well.
Meanwhile, the proposed model does not represent the real world because it is made up of minimum simple assumptions. In order to propose a model that is close to reality, more physical axioms should be reflected in the model. However, the advantage of deep learning algorithms is the fact that they are flexible in constructing the architecture. For example, in order for the operator to satisfy covariance, the weight sharing condition should be enforced in the proposed deep learning architecture[10]. In addition, the activation function must be carefully selected to represent the CNN in covariant form. For the case of G-CNN, activation function satisfies covariance when a pointwise function is used for activation. In this way, it can be expected that the more physical requirements are mapped to the components that make up deep learning architecture, the closer the model will be to reality.

### v8

@@ III A Deep Learning Algorithm for a Holographic System
Here, white represents a living cell and black represents a dead cell. During the inference process, the well-known glider pattern from the game of life was used as the input. However, it can be confirmed that the output after applying is the same as the output after applying , even with different inputs. Therefore, the proposed algorithm can learn a function that decrypts the information from one space to another by leveraging commutativity. Because deep learning is a non-linear transform, this model could include non-trivial interactions if more physical constraints are added to the proposed model.
IV Discussion
In the proposed model, it was essential to assume that a time evolution law exists for a many-body system, as in cellular automata. Indeed, this rule is not as simple as block CA and can not be expressed in a mathematically closed form. Therefore, if this rule exists, how to find it is a difficult problem. But again, one might also expect this rule to be discovered through a deep learning algorithm. First, we have a way to test a system in the standard model. Therefore, if one collects sufficient datasets via multiple experiments, the evolution law can be learned. In fact, if we assume that random data really evolves according to the block CA rule as an example in the previous section, it has been confirmed that the block CA rule can be learned through an appropriate deep learning architecture and numerous random dataset. Moreover, it can be confirmed that learning is possible in reverse time as well.
However, as previously noted, the curvature of the bulk causes the time evolution operator and the evolution operator along the holographic direction to form a closely related pair. In fact, while the time evolution operators should be identical on the two hypersurfaces, the operator along the holographic direction must be capable of exhibiting a time delay. Such a time evolution operator is unlikely to be a mere block CA. Nonetheless, identifying even a simple example of a paired set of evolution operators that satisfies these conditions could be a valuable step forward, offering a promising direction for advancing the current research.
Meanwhile, the proposed model does not represent the real world because it is made up of minimum simple assumptions. In order to propose a model that is close to reality, more physical principles should be reflected in the model. However, the advantage of deep learning algorithms is the fact that they are flexible in constructing the architecture. For example, in order for the operator to satisfy general covariance also, the weight sharing condition should be enforced in the proposed deep learning architecture[14]. In addition, the activation function must be carefully selected to represent the CNN in covariant form. For the case of G-CNN, activation function satisfies covariance when a pointwise function is used for activation. In this way, it can be expected that mapping more physical requirements to the components of the deep learning architecture will bring the model closer to reality.